OpenStudy (anonymous):

help me derivatives again, f(x)= (2x^2 + 5)^3 (4x-1) diff calculus. can you show me the proper way to solve to easy to understand. please

OpenStudy (jhannybean):

\[f(x) =(2x^2+5)^3(4x-1)\]You could... use the product rule to find the derivative, perhaps. \[d(f(x)) = f'g +g'f\]

OpenStudy (anonymous):

so how can we solve that?

OpenStudy (jhannybean):

\[d(f(x)) = [(2x^2+5)^3]'\cdot (4x-1) +(4x-1)'\cdot (2x^2+5)^3\]

OpenStudy (jhannybean):

Can you take it from there?

OpenStudy (anonymous):

ill get my reading glass

OpenStudy (jhannybean):

Hows it going? :)

OpenStudy (anonymous):

how can we solve that?

OpenStudy (jhannybean):

Do you know how to use the chain rule?

OpenStudy (anonymous):

no

OpenStudy (jhannybean):

\[d(x^n) = nx^{n-1}\]This?

OpenStudy (anonymous):

no sorry. :((

OpenStudy (jhannybean):

Chain rule works like this: \[d(f(g(x)) = [f(x)]' \cdot [g(x)]'\]So.. you have a function within a function. \[f(x) = (2x^2+5)^3\]\[g(x) = 2x^2+5\]Are you with me so far?

OpenStudy (anonymous):

yes im with you :)

OpenStudy (anonymous):

so what our the final answr?

OpenStudy (jhannybean):

Well, I am guiding you through to find it :)

OpenStudy (anonymous):

but its hard

OpenStudy (jhannybean):

Yes, calculus is quite difficult.

OpenStudy (jhannybean):

Stay with me here though, you will learn how to do this :)

OpenStudy (anonymous):

yes mam/ :))

OpenStudy (jhannybean):

So first we take the derivative of our "outer" function, \(f(x)\). In order to do that, we apply the power rule, \(\large \frac{d}{dx}(x^n) = n\cdot x^{n-1}\). Basically, we bring down the power of the function, and reduce it by 1. So \[\frac{d}{dx} (f(x)) = 3(2x^2+5)^{3-1} = 3(2x^2+5)^2\]

OpenStudy (jhannybean):

Did you understand that?

OpenStudy (anonymous):

ah i try to get it

OpenStudy (anonymous):

so our answr should be 3(2x^2+5)^2

OpenStudy (jhannybean):

That is not our answer yet, that is just a part of it.

OpenStudy (anonymous):

ahh sorrry mam

OpenStudy (jhannybean):

Now we take our "inner" function, and we also the power rule to it. Then we multiply it to the derivative of our function, \(f(x)\)\[\frac{d}{dx}(g(x)) = (2x^2+5)' = 4x\]

OpenStudy (jhannybean):

So for the first portion, we will have \[\frac{d}{dx}(2x^2+5)^3) = 3(2x^2+5)^2\cdot4x = 12x(2x^2+5)^2\]

OpenStudy (jhannybean):

Are you undertanding any of this?

OpenStudy (anonymous):

yes mam

OpenStudy (anonymous):

so if we get the power role? mam what will happen to our solving?

OpenStudy (jhannybean):

I don't understand your question?

OpenStudy (anonymous):

what will happen to our solving? if we get the power rule

OpenStudy (jhannybean):

The power rule is the method we use to find derivatives, it helps make functions easier to combine and solve for.

OpenStudy (jhannybean):

But anyway, So we started with this: \[d(f(x)) = \color{red}{[(2x^2+5)^3]'}\cdot (4x-1) +(4x-1)'\cdot (2x^2+5)^3\]And that highlighted portion we found, turned into this:So we started with this: \[d(f(x)) = [\color{red}{12x(2x^2+5)^2}\cdot (4x-1) +(4x-1)'\cdot (2x^2+5)^3\]

OpenStudy (anonymous):

yes mam

OpenStudy (jhannybean):

Now we do the same thing with our \(g(x)\) function, \((4x-1)\).\[\frac{d}{dx} (4x-1) = \frac{d}{dx}(4x) = 4\]

OpenStudy (anonymous):

yes mam

OpenStudy (jhannybean):

\[d(f(x)) = [12x(2x^2+5)^2\cdot (4x-1) +\color{red}{(4x-1)'}\cdot (2x^2+5)^3\]\[d(f(x)) = 12x(2x^2+5)^2\cdot (4x-1) +\color{red}4\cdot (2x^2+5)^3\]

OpenStudy (jhannybean):

So now, we see that \((2x^2+5)^2\) is common throughout the function, we can factor that out.\[d(f(x)) = 12x(2x^2+5)^2\cdot (4x-1) +4\cdot (2x^2+5)^3\]\[f'(x)=(2x^2+5)^2[12x(4x-1)+4(2x^2+5)]\]

OpenStudy (jhannybean):

Now you can simplify what is in the brackets and see if anything cancels out, or combines, or you can leave it that way.

OpenStudy (anonymous):

yes mam

OpenStudy (jhannybean):

before doing any more derivatives, read through your calculus book to learn how to find them. That will come in handy.

OpenStudy (jhannybean):

Good luck :)

OpenStudy (anonymous):

mam

OpenStudy (anonymous):

last question!

OpenStudy (jhannybean):

yes?

OpenStudy (anonymous):

how can i cet the final answr?

OpenStudy (jhannybean):

Simplify what is inside the brackets.

OpenStudy (anonymous):

how>?

OpenStudy (jhannybean):

I have worked most of the problem for you, I am sure you know how to combine like terms :)

OpenStudy (anonymous):

mam

OpenStudy (jhannybean):

I have got to go now, I hope you practice these more :)

OpenStudy (anonymous):

thank you, :)) You know I was thinking, I don't want you to be my math teacher. I don't want you to teach me. Because I like you so much.

OpenStudy (jhannybean):

Good luck :)

OpenStudy (anonymous):

just good luck? @Jhannybean